From point A, 100m due south of a tower, the angle of elevation of the top of the tower is 40 degrees. From point B due east of the tower, the angle of elevation of the top of the tower is 20 degrees. How far apart are points A and B?
1
Expert's answer
2020-03-27T17:46:52-0400

"OC=OA\tan40^0" ,"OB=\frac{OC}{\tan20^0}" , "AB=\sqrt {OA^2+OB^2}=OA\sqrt{1+\frac{\tan^240^0}{\tan^220^0}}\approx252m"

if "OA=100m"